# Incorporating Arbitrary Matrix Group Equivariance into KANs

### v2

@@ 4.3 Equivariant linear weights
The output feature $v_{go}\in U_{go}$ is obtained by a linear combination of the post-activation value $v_{m}\in U_{m}$. Let $U_{i}=\mathbb{R}^{d_{i}}$ and $U_{go}=\mathbb{R}^{d_{go}}$, then Eq. 15 indicates that $U_{m}=\mathbb{R}^{(G+k+1)d_{i}}$. The linear weight matrix $W\in\mathbb{R}^{d_{go}\times(G+k+1)d_{i}}$ can be partitioned as $W=[W_{0}\ W_{1}\ \dots\ W_{G+k}]$, where $W_{b}\in\mathbb{R}^{d_{go}\times d_{i}}$. Combining with Eq. 13, we have:
@@
where $\rho_{go}$ is the group representation of $U_{go}$. Using the structure of $U_{m}$ shown in Eq. 15 and applying the rules from Eq. 4, we can derive that $\rho_{m}(g)=\bigoplus_{b=0}^{G+k}\rho_{i}(g)$, where $\rho_{i}$ is the group representation of $U_{i}$. Therefore, from Eq. 13, we have $\rho_{m}(g)v_{m}=\bigoplus_{b=0}^{G+k}\rho_{i}(g)v_{m,b}$. Then Eq. 18 can be written as:
@@
where $\rho_{go,i}=\rho_{go}\otimes\rho_{i}^{*}$ is the group representation of $U_{go}\otimes U_{i}^{*}$. We use the same method as EMLP [17] to solve for the equivariant basis $Q$ and equivariant projector $P=QQ^{\top}$ of $\mathrm{vec}(W_{b})$. Similar to the transition from Eq. 6 to Eq. 7, we decompose the group representation $\rho_{go,i}(g)$ into discrete and infinitesimal generators to obtain the following constraint:
@@
We summarize the architecture of the EKAN layer in Fig. 2 (Right). In this example, the EKAN layer is equivariant with respect to a 2-dimensional matrix group $\widetilde{G}$ (such as the $\mathrm{SO}(2)$ group). The user specifies the input space $U_{i}=T_{0}\oplus T_{1}$ (where we abbreviate $T(p,0)=V^{p}$ as $T_{p}$), which represents a scalar space and a vector space, and specifies the output space $U_{o}=T_{2}$, which represents a matrix space. Then, the gated input space $U_{gi}=T_{0}\oplus T_{1}\oplus T_{0}$ and the gated output space $U_{go}=T_{2}\oplus T_{0}$ each add a gate scalar to the vector space $T_{1}$ and the matrix space $T_{2}$. The basis functions are applied to the gate scalars of $T_{0}$ (itself) and $T_{1}$, which are then multiplied by the original terms to obtain the post-activation space $U_{m}=3U_{i}$. The linear weights $W\in\mathbb{R}^{5\times 9}$ between $U_{m}$ and $U_{go}$ are within the subspace determined by Sec. 4.3 to ensure equivariance.
4.4 Grid update
Similar to KANs [37], EKAN updates grids based on the input activations. At the same time, the linear weights should also be updated in order to keep the output features unchanged. Let the post-activation values of the grids before and after the update be denoted as $V_{m},V_{m}^{\prime}\in\mathbb{R}^{N\times(G+k+1)d_{i}}$, where $N$ is the number of samples. We first project the linear weight blocks into the equivariant subspace as $\mathrm{vec}(\widetilde{W}_{b})=P\mathrm{vec}(W_{b})$ and compute the output activations $V_{go}=V_{m}\widetilde{W}^{\top}=\sum_{b=0}^{G+k}V_{m,b}\widetilde{W}_{b}^{\top}$, where $P$ is the equivariant projector obtained from Sec. 4.3. Then, the updated equivariant linear weights $\widetilde{W}^{\prime}$ should satisfy $V_{go}=V_{m}^{\prime}\widetilde{W}^{\prime\top}$, and we have $\widetilde{W}^{\prime}=V_{go}^{\top}(V_{m}^{\prime\top})^{\dagger}$. We finally restore the updated linear weight blocks $\mathrm{vec}(W_{b}^{\prime})=P^{\dagger}\mathrm{vec}(\widetilde{W}_{b}^{\prime})$.
@@ 5 EKAN architecture
However, the actual input/output features of the EKAN layer stack lie in $U_{gi}$/$U_{go}$. Therefore, we need to add extensions to align the network with the dataset. First, the gate scalars of the actual output feature are directly dropped to obtain the final output label of EKAN, which resides in $U_{o}$. Then, we add a lift layer before the first layer to preprocess the raw input feature of EKAN, which is essentially an equivariant linear layer between $U_{i}$ and $U_{gi}$ (see Supplementary Material for more details).
We summarize the space structures and network architecture of EKAN in Fig. 3. The space structures of EKAN can be analogous to the dimensions of a conventional network. The user specifies the latent space $U_{l}$ of EKAN, which corresponds to specifying the hidden dimension in a conventional network. The input space $U_{i}$ and the output space $U_{o}$ are determined by the dataset, similar to how the input and output dimensions are defined in a conventional network. In the concrete example, the feature space and label space of the dataset are $U_{i}=T_{0}\oplus T_{1}$ and $U_{o}=T_{2}$, respectively. After passing through the lift layer, a new gate scalar is added to the raw input feature for $T_{1}$, resulting in the actual input feature $U_{gi}=T_{0}\oplus T_{1}\oplus T_{0}$ for the first EKAN layer. The gate scalar in the actual output space $U_{go}=T_{2}\oplus T_{0}$ of the last EKAN layer is dropped to obtain the final output label.
@@ 6 Experiments
In this section, we evaluate the performance of EKAN on regression and classification tasks with known symmetries. Compared with similarly general architectures such as MLPs, KANs, and EMLP, EKAN achieves lower test loss and higher test accuracy with smaller datasets or fewer parameters.
6.1 Particle scattering
In electron-muon scattering, we can observe the four-momenta of the incoming electron, incoming muon, outgoing electron, and outgoing muon, denoted as $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\in\mathbb{R}^{4}$, respectively. We aim to predict the matrix element, which is proportional to the cross-section [17]:
$|\mathcal{M}|^{2}\propto[p^{\mu}\tilde{p}^{\nu}-(p^{\alpha}\tilde{p}_{\alpha}-p^{\alpha}p_{\alpha})g^{\mu\nu}][q_{\mu}\tilde{q}_{\nu}-(q^{\alpha}\tilde{q}_{\alpha}-q^{\alpha}q_{\alpha})g_{\mu\nu}].$ (23)
According to Einstein’s summation convention, in a monomial, if an index appears once as a superscript and once as a subscript, it indicates summation over that index. The metric tensor is given by $g_{\mu\nu}=g^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$, and $a_{\mu}=g_{\mu\nu}a^{\nu}=(a^{0},-a^{1},-a^{2},-a^{3})$. The matrix element is invariant under Lorentz transformations. In other words, this task exhibits $\mathrm{O}(1,3)$ invariance (see Supplementary Material for more details), with the feature space $U_{i}=4T_{1}$ and the label space $U_{o}=T_{0}$.
We embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN. Models are evaluated on synthetic datasets with different training set sizes, which are generated by sampling $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\sim\mathcal{N}(0,\frac{1}{4^{2}})$. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. Although the lift layer increases the parameter overhead, we set the width of the middle layer in EKAN to $n_{1}=1000$ (shape as $[16,1000,1]$, and the software will automatically calculate the appropriate feature space structure based on the user-specified dimension), and set the width of the middle layer in KAN to $n_{1}=3840$ (shape as $[16,3840,1]$) to keep the parameter count similar. Both EMLP and MLP have the depth of $L=4$ and the middle layer width of $n_{1}=n_{2}=n_{3}=384$ (shape as $[16,384,384,384,1]$). In these settings, EKAN ($435$k) has fewer parameters than EMLP ($450$k) and KAN ($461$k). We provide more implementation details in Supplementary Material.
@@ 6.1 Particle scattering
We repeat experiments with three different random seeds and report the mean $\pm$ std of the test MSE in Tab. 1. The results of EMLP and MLP come from the original paper [17] under the same settings. Although EMLP performs better than non-equivariant models, our EKAN with different group equivariance further surpasses it comprehensively, especially showing an orders-of-magnitude advantage on large datasets (training set size $\geq 10^{3}$). Moreover, our EKAN with just $10^{3}$ training samples achieves approximately $10\%$ of the test MSE of EMLP with $10^{4}$ training samples.
6.2 Three-body problem
The study of the three-body problem on a plane [23] focuses on the motion of three particles, with their center of mass at the origin, under the influence of gravity. Their trajectories are chaotic and cannot be described by an analytical solution. Specifically, we observe the motion states of three particles over the past four time steps, denoted as $\{q_{i1},p_{i1},q_{i2},p_{i2},q_{i3},p_{i3}\}_{i=t-4}^{t-1}$, and predict their motion states at time $t$, denoted as $\{q_{t1},p_{t1},q_{t2},p_{t2},q_{t3},p_{t3}\}$. Here, $q_{ij}\in\mathbb{R}^{2}$ and $p_{ij}\in\mathbb{R}^{2}$ indicate the position and momentum coordinates of the $j$-th particle at time $i$, respectively. The dataset contains $30$k training samples and $30$k test samples. When the input motion states are simultaneously rotated by a certain angle or reflected along a specific axis, the output motion states should undergo the same transformation. Therefore, this task has $\mathrm{O}(2)$ equivariance (see Supplementary Material for more details), with the feature space $U_{i}=4\times 6T_{1}=24T_{1}$ and the label space $U_{o}=6T_{1}$.
We embed the group $\mathrm{O}(2)$ and its subgroup $\mathrm{SO}(2)$ equivariance into EKAN and EMLP. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$, while both EMLP and MLP have the depth of $L=4$. The number of parameters is controlled by adjusting the middle layer width $N$ for comparison (the shape of EKAN and KAN is $[48,N,12]$, while the shape of EMLP and MLP is $[48,N,N,N,12]$). More implementation details can be found in Supplementary Material.
@@ 6.2 Three-body problem
The mean $\pm$ std of the test MSE over three runs with different random seeds are reported in Tab. 2. Our EKAN-$\mathrm{SO}(2)$ and EKAN-$\mathrm{O}(2)$ consistently outperform baseline models with the same number of parameters, often by orders of magnitude. Notably, our EKAN with $10^{4.5}$ parameters achieves comparable or even lower test MSE than baseline models with $10^{5.5}$ parameters, saving $90\%$ of the parameter overhead.
6.3 Top quark tagging
The research on top quark tagging [29] involves classifying hadronic tops from the QCD background. In particle collision experiments, top quark decays or other events produce several jet constituents. We observe the four-momenta $p_{1}^{\mu},p_{2}^{\mu},p_{3}^{\mu}\in\mathbb{R}^{4}$ of the three jet constituents with the highest transverse momentum $p_{T}$, and predict the event label ($1$ for top, 00 for QCD). The category of the event will not change when all jet constituents undergo the same Lorentz transformation. Consequently, this task possesses $\mathrm{O}(1,3)$ invariance (see Supplementary Material for more details), with the feature space $U_{i}=3T_{1}$ and the label space $U_{o}=T_{0}$.
Some previous works specifically focus on top quark tagging [39, 4, 21] have achieved good results. However, EKAN aims to provide a general framework for all symmetric scientific problems, so methods targeted at downstream tasks are not our primary benchmark for comparison. Combining EKAN with task-specific models remains a direction for future research.
Similar to particle scattering, we embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN and EMLP. We sample training sets of different sizes from the original dataset for evaluation. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. We set the width of the middle layer in EKAN to $n_{1}=200$ (shape as $[12,200,1]$) and the width of the middle layer in KAN to $n_{1}=384$ (shape as $[12,384,1]$) to control the number of parameters. Both EMLP and MLP have the depth of $L=4$ and the middle layer width $n_{1}=n_{2}=n_{3}=200$ (shape as $[12,200,200,200,1]$). We apply the sigmoid function to the model’s output and use BCE as the loss function for binary classification. More implementation details are provided in Supplementary Material.
@@ 6.3 Top quark tagging
[…]
We report the mean $\pm$ std of the test accuracy over three runs with different random seeds, as well as the number of parameters of the models in Tab. 3. Since we have not observed all the jet constituents, the relationship between the labels and input features cannot be accurately expressed as an explicit function. In this case of non-symbolic formula representation, KAN cannot achieve higher accuracy with fewer parameters than MLP as expected. On the other hand, our EKAN achieves comparable results with EMLP using approximately $26\%$ of the parameters, improving test accuracy by $0.23\%\sim 6.44\%$ on small datasets (training set size $<10^{4}$), while decreasing by $0.19\%$ on large datasets (training set size $=10^{4}$).

### v3

@@ 4.3 Equivariant linear weights
The output feature $v_{go}\in U_{go}$ is obtained by a linear combination of the post-activation value $v_{m}\in U_{m}$. Let $U_{i}=\mathbb{R}^{d_{i}}$ and $U_{go}=\mathbb{R}^{d_{go}}$, then Equation 13 indicates that $U_{m}=\mathbb{R}^{(G+k+1)d_{i}}$. The linear weight matrix $W\in\mathbb{R}^{d_{go}\times(G+k+1)d_{i}}$ can be partitioned as $W=[W_{0}\ W_{1}\ \dots\ W_{G+k}]$, where $W_{b}\in\mathbb{R}^{d_{go}\times d_{i}}$. Combining with Equation 11, we have:
@@
where $\rho_{go}$ is the group representation of $U_{go}$. Using the structure of $U_{m}$ shown in Equation 13 and applying the rules from Equation 4, we can derive that $\rho_{m}(g)=\bigoplus_{b=0}^{G+k}\rho_{i}(g)$, where $\rho_{i}$ is the group representation of $U_{i}$. Therefore, from Equation 11, we have $\rho_{m}(g)v_{m}=\bigoplus_{b=0}^{G+k}\rho_{i}(g)v_{m,b}$. Then Equation 16 can be written as:
@@
where $\rho_{go,i}=\rho_{go}\otimes\rho_{i}^{*}$ is the group representation of $U_{go}\otimes U_{i}^{*}$. We use the same method as EMLP (Finzi et al., 2021) to solve for the equivariant basis $Q$ and equivariant projector $P=QQ^{\top}$ of $\mathrm{vec}(W_{b})$. Similar to the transition from Equation 27 to Equation 28 (in Appendix B), we decompose the group representation $\rho_{go,i}(g)$ into discrete and infinitesimal generators to obtain the following constraint:
@@
We summarize the architecture of the EKAN layer in Figure 2 (Right). In this example, the EKAN layer is equivariant with respect to a 2-dimensional matrix group $\widetilde{G}$ (such as the $\mathrm{SO}(2)$ group). The user specifies the input space $U_{i}=T_{0}\oplus T_{1}$ (where we abbreviate $T(p,0)=V^{p}$ as $T_{p}$), which represents a scalar space and a vector space, and specifies the output space $U_{o}=T_{2}$, which represents a matrix space. Then, the gated input space $U_{gi}=T_{0}\oplus T_{1}\oplus T_{0}$ and the gated output space $U_{go}=T_{2}\oplus T_{0}$ each add a gate scalar to the vector space $T_{1}$ and the matrix space $T_{2}$. The basis functions are applied to the gate scalars of $T_{0}$ (itself) and $T_{1}$, which are then multiplied by the original terms to obtain the post-activation space $U_{m}=3U_{i}$. The linear weights $W\in\mathbb{R}^{5\times 9}$ between $U_{m}$ and $U_{go}$ are within the subspace determined by Section 4.3 to ensure equivariance. Similar to KANs (Liu et al., 2024b), EKAN updates grids based on the input activations, which we discuss in detail in Appendix D.
@@ 5 EKAN architecture
However, the actual input/output features of the EKAN layer stack lie in $U_{gi}$/$U_{go}$. Therefore, we need to add extensions to align the network with the dataset. First, the gate scalars of the actual output feature are directly dropped to obtain the final output label of EKAN, which resides in $U_{o}$. Then, we add a lift layer before the first layer to preprocess the raw input feature of EKAN, which is essentially an equivariant linear layer between $U_{i}$ and $U_{gi}$ (see Section 4.3).
We summarize the space structures and network architecture of EKAN in Figure 3. The space structures of EKAN can be analogous to the dimensions of a conventional network. The user specifies the latent space $U_{l}$ of EKAN, which corresponds to specifying the hidden dimension in a conventional network. The input space $U_{i}$ and the output space $U_{o}$ are determined by the dataset, similar to how the input and output dimensions are defined in a conventional network. In the concrete example, the feature space and label space of the dataset are $U_{i}=T_{0}\oplus T_{1}$ and $U_{o}=T_{2}$, respectively. After passing through the lift layer, a new gate scalar is added to the raw input feature for $T_{1}$, resulting in the actual input feature $U_{gi}=T_{0}\oplus T_{1}\oplus T_{0}$ for the first EKAN layer. The gate scalar in the actual output space $U_{go}=T_{2}\oplus T_{0}$ of the last EKAN layer is dropped to obtain the final output label.
@@ 6 Experiments
MLP + augmentation $(1.90\pm 0.22)\times 10^{-1}$ $(3.97\pm 0.52)\times 10^{-2}$ $(1.25\pm 0.08)\times 10^{-2}$ $(1.08\pm 0.03)\times 10^{-2}$ $(1.34\pm 0.22)\times 10^{-2}$
@@
KAN + augmentation $(5.98\pm 0.67)\times 10^{-1}$ $(4.72\pm 1.52)\times 10^{-1}$ $(9.07\pm 2.36)\times 10^{-2}$ $(5.61\pm 0.70)\times 10^{-2}$ $(1.58\pm 0.06)\times 10^{-1}$
@@
MLP + augmentation $(9.43\pm 0.88)\times 10^{-3}$ $(9.65\pm 0.54)\times 10^{-3}$ $(1.01\pm 0.09)\times 10^{-2}$ $(9.89\pm 0.65)\times 10^{-3}$ $(9.91\pm 0.46)\times 10^{-3}$
@@
KAN + augmentation $(7.94\pm 0.15)\times 10^{-3}$ $(7.15\pm 0.24)\times 10^{-3}$ $(6.91\pm 0.34)\times 10^{-3}$ $(6.88\pm 0.06)\times 10^{-3}$ $(6.76\pm 0.12)\times 10^{-3}$
@@
In this section, we evaluate the performance of EKAN on regression and classification tasks with known symmetries. Compared with general models such as MLPs, KANs, and equivariant architectures like EMLP (Finzi et al., 2021) and CGENN (Ruhe et al., 2023), EKAN achieves lower test loss and higher test accuracy with smaller datasets or fewer parameters.
6.1 Particle scattering
In electron-muon scattering, we can observe the four-momenta of the incoming electron, incoming muon, outgoing electron, and outgoing muon, denoted as $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\in\mathbb{R}^{4}$, respectively. We aim to predict the matrix element, which is proportional to the cross-section (Finzi et al., 2021): $|\mathcal{M}|^{2}\propto[p^{\mu}\tilde{p}^{\nu}-(p^{\alpha}\tilde{p}_{\alpha}-p^{\alpha}p_{\alpha})g^{\mu\nu}][q_{\mu}\tilde{q}_{\nu}-(q^{\alpha}\tilde{q}_{\alpha}-q^{\alpha}q_{\alpha})g_{\mu\nu}]$. According to Einstein’s summation convention, in a monomial, if an index appears once as a superscript and once as a subscript, it indicates summation over that index. The metric tensor is given by $g_{\mu\nu}=g^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$, and $a_{\mu}=g_{\mu\nu}a^{\nu}=(a^{0},-a^{1},-a^{2},-a^{3})$. The matrix element is invariant under Lorentz transformations. In other words, this task exhibits $\mathrm{O}(1,3)$ invariance (see Section A.2 for more details), with the feature space $U_{i}=4T_{1}$ and the label space $U_{o}=T_{0}$.
We embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN. Models are evaluated on synthetic datasets with different training set sizes, which are generated by sampling $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\sim\mathcal{N}(0,\frac{1}{4^{2}})$. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. Although the lift layer increases the parameter overhead, we set the width of the middle layer in EKAN to $n_{1}=1000$ (shape as $[16,1000,1]$, and the software will automatically calculate the appropriate feature space structure based on the user-specified dimension), and set the width of the middle layer in KAN to $n_{1}=3840$ (shape as $[16,3840,1]$) to keep the parameter count similar. Both EMLP and MLP have the depth of $L=4$ and the middle layer width of $n_{1}=n_{2}=n_{3}=384$ (shape as $[16,384,384,384,1]$). In these settings, EKAN ($435$k) has fewer parameters than EMLP ($450$k) and KAN ($461$k). We provide more implementation details in Section E.1.
We repeat experiments with three different random seeds and report the mean $\pm$ std of the test MSE in Table 1. The results of EMLP and MLP come from the original paper (Finzi et al., 2021) under the same settings. Although EMLP performs better than non-equivariant models, our EKAN with different group equivariance further surpasses it comprehensively, especially showing an orders-of-magnitude advantage on large datasets (training set size $\geq 10^{3}$). Moreover, our EKAN with just $10^{3}$ training samples achieves approximately $10\%$ of the test MSE of EMLP with $10^{4}$ training samples.
6.2 Three-body problem
[…]
@@ 6.2 Three-body problem
The study of the three-body problem on a plane (Greydanus et al., 2019) focuses on the motion of three particles, with their center of mass at the origin, under the influence of gravity. Their trajectories are chaotic and cannot be described by an analytical solution. Specifically, we observe the motion states of three particles over the past four time steps, denoted as $\{q_{i1},p_{i1},q_{i2},p_{i2},q_{i3},p_{i3}\}_{i=t-4}^{t-1}$, and predict their motion states at time $t$, denoted as $\{q_{t1},p_{t1},q_{t2},p_{t2},q_{t3},p_{t3}\}$. Here, $q_{ij}\in\mathbb{R}^{2}$ and $p_{ij}\in\mathbb{R}^{2}$ indicate the position and momentum coordinates of the $j$-th particle at time $i$, respectively. The dataset contains $30$k training samples and $30$k test samples. When the input motion states are simultaneously rotated by a certain angle or reflected along a specific axis, the output motion states should undergo the same transformation. Therefore, this task has $\mathrm{O}(2)$ equivariance (see Section A.1 for more details), with the feature space $U_{i}=4\times 6T_{1}=24T_{1}$ and the label space $U_{o}=6T_{1}$.
We embed the group $\mathrm{O}(2)$ and its subgroup $\mathrm{SO}(2)$ equivariance into EKAN and EMLP. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$, while both EMLP and MLP have the depth of $L=4$. The number of parameters is controlled by adjusting the middle layer width $N$ for comparison (the shape of EKAN and KAN is $[48,N,12]$, while the shape of EMLP and MLP is $[48,N,N,N,12]$). More implementation details can be found in Section E.2.
The mean $\pm$ std of the test MSE over three runs with different random seeds are reported in Table 2. Our EKAN-$\mathrm{SO}(2)$ and EKAN-$\mathrm{O}(2)$ consistently outperform baseline models with the same number of parameters, often by orders of magnitude. Notably, our EKAN with $10^{4.5}$ parameters achieves comparable or even lower test MSE than baseline models with $10^{5.5}$ parameters, saving $90\%$ of the parameter overhead.
6.3 Top quark tagging
The research on top quark tagging (Kasieczka et al., 2019) involves classifying hadronic tops from the QCD background. In particle collision experiments, top quark decays or other events produce several jet constituents. We observe the four-momenta $p_{1}^{\mu},p_{2}^{\mu},p_{3}^{\mu}\in\mathbb{R}^{4}$ of the three jet constituents with the highest transverse momentum $p_{T}$, and predict the event label ($1$ for top, 00 for QCD). The category of the event will not change when all jet constituents undergo the same Lorentz transformation. Consequently, this task possesses $\mathrm{O}(1,3)$ invariance (see Section A.2 for more details), with the feature space $U_{i}=3T_{1}$ and the label space $U_{o}=T_{0}$.
Similar to particle scattering, we embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN and EMLP. Furthermore, we embed $\mathrm{O}(1,3)$ equivariance into CGENN (Ruhe et al., 2023). We sample training sets of different sizes from the original dataset for evaluation. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. We set the width of the middle layer in EKAN to $n_{1}=200$ (shape as $[12,200,1]$) and the width of the middle layer in KAN to $n_{1}=384$ (shape as $[12,384,1]$) to control the number of parameters. Both MLP, EMLP and CGENN have the depth of $L=4$ and the middle layer width $n_{1}=n_{2}=n_{3}=200$ (shape as $[12,200,200,200,1]$). We apply the sigmoid function to the model’s output and use BCE as the loss function for binary classification. More implementation details are provided in Section E.3.
We report the mean $\pm$ std of the test accuracy over three runs with different random seeds, as well as the number of parameters of the models in Table 3. Since we have not observed all the jet constituents, the relationship between the labels and input features cannot be accurately expressed as an explicit function. In this case of non-symbolic formula representation, KAN cannot achieve higher accuracy with fewer parameters than MLP as expected. On the other hand, our EKAN achieves comparable results with EMLP and CGENN using fewer than $40\%$ of the parameters, improving test accuracy by $0.23\%\sim 6.44\%$ on small datasets (training set size $<10^{4}$), while decreasing by $0.19\%$ on large datasets (training set size $=10^{4}$).